OpenStudy (anonymous):

Suppose that sec(B) = - 8⁄3 and that 180o < B < 270o. Find sin(B).

OpenStudy (anonymous):

Suppose that sec(B) = - 12⁄5 and that 180o < B < 270o. Find sin(B)

hartnn (hartnn):

since sec B=1/cos B so 1st find cos B. Then use \(sin^2 B+cos ^2 B=1\) to find sin B

OpenStudy (anonymous):

thanks hartnn

hartnn (hartnn):

welcome. :) tell me what answer u got,though.

OpenStudy (anonymous):

k hang on

OpenStudy (anonymous):

@ hartnn this is what i get sin b =13/12

hartnn (hartnn):

nopes,how did u get that? whats cos B u got?

OpenStudy (anonymous):

cos b = 5/-12

OpenStudy (anonymous):

can u work it and i c where i went wrong? please

hartnn (hartnn):

u might have got sin^2 B = 169/144 right ??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

then i found the root...is that not what I was supposed to do?

hartnn (hartnn):

nopes, even thats not correct.....let me work it out.. cos B =-5/12 sin^2 B= 1- 25/144 sin^2 B = 119/144 clear till here ???

OpenStudy (anonymous):

how do u get 119?

hartnn (hartnn):

144-25 \(1-\frac{25}{144}=\frac{144-25}{144}=\frac{119}{144}\)

OpenStudy (anonymous):

okay gat you..

OpenStudy (anonymous):

continue please

hartnn (hartnn):

now as 180o < B < 270o. B is in 3rd Quadrant, in this quadrant, sine is negative. so sin B=-\(\sqrt{\frac{119}{144}}=-\frac{\sqrt{119}}{12}\) ok ?

OpenStudy (anonymous):

wow!! you're a genius thanks alot

OpenStudy (anonymous):

I have some more please don go away..

hartnn (hartnn):

welcome :) ask if any more doubts.

OpenStudy (anonymous):

can you make up on so we can try it again and c if i get the same answer as yours?

hartnn (hartnn):

ofcourse,try the one u have asked in the question.....

OpenStudy (anonymous):

cool

hartnn (hartnn):

1-9/64=55/64 so \(sin B=-\sqrt{\frac{55}{64}}=-\frac{\sqrt{55}}{8}\)